OpenStudy (anonymous):

Use part I of the Fundamental Theorem of Calculus to find the derivative of h(x) = \int_{-1}^{\sin(x)} (\cos(t^2)+t)\; dt

OpenStudy (ash2326):

\[h(x) = \int_{-1}^{\sin(x)} (\cos(t^2)+t)\; dt\]

OpenStudy (anonymous):

-2cosx(cos(sin^2x))+sinx my answer. Thanks ash for puting up the right format. I can't figureout how my answer is wrong though

OpenStudy (ash2326):

You put t= sin x and multiply the expression by the derivative of sin x

OpenStudy (anonymous):

that's what I did but somehow I got the wrong answer. could you look at my answer and tell me why it's wrong?

OpenStudy (ash2326):

t= sin x, as the lower limit is constant it'd be 0 \[\{\frac{d}{dx} (\sin x) \}\times ( \cos(\sin^2 x))+\sin x )-0\] \[\cos x ( \cos(\sin^2 x))+\sin x )\]

OpenStudy (ash2326):

\[ \cos x(\cos(\sin^2 x)+\sin x )\]

OpenStudy (ash2326):

Did you get it?

OpenStudy (anonymous):

I see where I went wrong thanks!